OpenStudy (anonymous):

Solve it with the Substitution method 3x+5y=-1 & y=2x-8

OpenStudy (anonymous):

I will work this one for you but check one of my youtube videos. Forgive the musical indulgence. http://www.youtube.com/watch?v=pNRGQV93yoA

OpenStudy (anonymous):

Write 3x + 5y = -1 (a) y = 2x - 8 (b) Now substitute the value of y in (b) into (a) Giving 3x + 5(2x- 8) = -1 Expand and simplify to find the value of x 3x +10x - 40 = -1 13x = 39 Therefore x = 3 Now substitute this value of x into (b) y = 2(3) - 8 Therefore y = 6 - 8 = -2 Answer = x is 3 and y is -2